OpenStudy (anonymous):

a rectangle is six times as long as it is wide. determine the ratio of its area to it's perimeter, in simplest form, if its width is w

OpenStudy (anonymous):

l*w = Area = 6w*w perimeter = 2w+2l = 2w +2*6w = 14w Area/perimeter = ...?

OpenStudy (anonymous):

Um I'm not quite sure